# Optimal Quoting under Adverse Selection and Price Reading

### v1

@@ 2.1 Modeling Framework and Notation
  • Adverse selection. For each $(n,k)\in\mathcal{N}\times\mathcal{K}$, the function $\tilde{\zeta}^{n,k}:\mathbb{R}\to\mathbb{R}_{+}$ is non-decreasing, and satisfies $\lim_{\delta\to-\infty}\tilde{\zeta}^{n,k}(\delta)=0.$ This function captures the impact of informed trading on prices. It is tier-specific (and possibly size-specific), allowing the market maker to account for varying levels of expected informativeness across clients (and transaction sizes). In particular, when $\tilde{\zeta}^{n,k}$ is a rapidly decreasing function, executions at prices that deviate significantly from the reference price in the “wrong” direction – that is, at prices seemingly unfavorable to the client from the market maker’s perspective – can have a strong impact on the reference price.
@@ 3.3.3 Optimal Quotes
$\displaystyle=$ $\displaystyle\delta^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}(\delta^{n,k,b*}(q))}\left(D^{k}_{+}f(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,b*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}\right)+o(\varepsilon)$
@@
$\displaystyle\mathfrak{d}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\delta^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}(\delta^{n,k,a*}(q))}\left(D^{k}_{-}f(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}\right$.
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.-\frac{(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,a*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}\right)+o(\varepsilon)$
These expressions highlight that the influence of informational risk on optimal quotes can be decomposed into two components. The first component arises from the terms $D^{k}_{\pm}f$, which are finite differences of the function $f$. By definition, these terms reflect the marginal impact of informational risk on the risk-adjusted expected PnL, assuming the market maker does not update their quotes. As such, they aggregate the effects of informational risk across all tiers and are identical for all $n$. The second component is represented by the terms $g^{n,k,b/a}$, which quantify the impact of informational risk on the greedy quotes derived from the baseline value function in the absence of informational risk. Unlike the $D^{k}_{\pm}f$ terms, the $g^{n,k,b/a}$ terms are specific to the considered tier and trade size.
@@ 4 Quadratic Hamiltonian Approximation
In the algorithmic market making literature, closed-form approximations of the value function and optimal quotes are frequently used. The standard approach (see [6]) consists in replacing the HJB equation associated with the control problem by an alternative HJB equation in which the Hamiltonian functions are approximated to second order. The resulting modified HJB equation admits a solution that is quadratic in the inventory variable, with coefficients computable through a Riccati equation and a linear equation. [^8] While we focus on the single-asset case in this article, it is worth emphasizing that this methodology extends naturally to multi-asset market making, thereby offering an effective remedy to the curse of dimensionality.
@@ 4.2 HJB Equation and Value Function
$A_{0}=\frac{-\rho+\sqrt{\rho^{2}+4\gamma\sigma^{2}F_{+}^{2,1}}}{4F_{+}^{2,1}}\qquad\text{and}\qquad B_{0}=-\frac{2A_{0}}{\rho+2A_{0}F_{+}^{2,1}}\left(A_{0}F_{-}^{2,2}+D_{-}^{1,1}\right),$
and
$A_{0}^{\prime}=-\frac{2A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\text{ and }B_{0}^{\prime}=-\frac{1}{\rho+2A_{0}F_{+}^{2,1}}\left(B_{0}\Sigma^{1,0}_{+}+A_{0}\Sigma^{3,1}_{-}+S-\frac{4A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\left(B_{0}F^{2,1}_{+}+2A_{0}F^{2,2}_{-}+F^{1,1}_{-}\right)\right).$
@@ 4.3 Optimal Quotes
If the greedy quotes in feedback form associated with $\widehat{\vartheta}_{0}$, i.e. in the case of no informational risk under the quadratic approximation, are denoted by
$\widehat{\delta}^{n,k,b*}(q):=\tilde{\delta}^{n,k,b*}\left(D^{k}_{+}\widehat{\vartheta}_{0}(q)\right)\quad\text{and}\quad\widehat{\delta}^{n,k,a*}(q):=\tilde{\delta}^{n,k,a*}\left(D^{k}_{-}\widehat{\vartheta}_{0}(q)\right),$
@@
$\displaystyle\widehat{\mathfrak{d}}^{n,k,b*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}\left(\widehat{\delta}^{n,k,b*}(q)\right)}\left(D^{k}_{+}\widehat{f}(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\widehat{\delta}^{n,k,b*}(q)\right)}\right$.
@@
$\displaystyle\widehat{\mathfrak{d}}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}\left(\widehat{\delta}^{n,k,a*}(q)\right)}\left(D^{k}_{-}\widehat{f}(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\widehat{\delta}^{n,k,a*}(q)\right)}\right$.
@@ 4.4 Analysis in the Symmetric Case
Given that $\widehat{f}$ is available in closed form, the above expressions are easier to handle than those in Section 3.3.3, which allows us to examine the underlying mechanisms in detail. For simplicity, we focus on a symmetric setting with identical bid- and ask-side functions, in the limiting case where $\rho\to 0$.
In this case indeed, we have $B_{\varepsilon}=B^{\prime}_{\varepsilon}=0$ and the first-order adjustment to the value function takes the form $\widehat{f}(q)=\frac{1}{2}\frac{\Sigma_{+}^{1,0}}{F_{+}^{2,1}}\,q^{2}$ (up to an additive constant) where [^10] We removed the $b$ and $a$ superscripts because of our assumption of symmetry.
@@
Let us first consider a tier where clients trade on strong signals and would execute in the direction of their signal at virtually any reasonable price given prevailing market conditions. Such behavior is plausible for clients acting on mid- or long-term signals. In this case, the slopes of the corresponding $\zeta$ functions are insufficiently steep to compensate the decrease of intensity functions, resulting in a reduction in the concavity of the value function. The mechanism is as follows: if the market maker holds a positive (resp. negative) inventory, then by skewing – through a decrease (resp. increase) in both bid and ask prices – they raise the absolute and relative probability of executing a sell (resp. buy) order that moves the inventory closer to zero and therefore the likelihood that adverse selection – whose dependence on quotes is mild – induces a price increase (resp. decrease), thereby artificially increasing the value of the remaining inventory. Although adverse selection reduces the constant term in the approximation of the value function, this mechanism tends to flatten it, thus lowering its concavity.
@@
The second component of the optimal quotes, which is tier-specific and relates to the local adjustments of greedy quotes, consists of two terms: one related to price reading and the other to adverse selection.
@@
Taken together, the two components of the impact of price reading on optimal quotes lead to a subtle trade-off. When the inventory is zero, the second component vanishes and only the first component remains: a widening of the bid-ask spread to account for informational risk. When the inventory is nonzero, the first component encourages stronger skewing to reduce the time spent in unbalanced positions, while the second component pushes in the opposite direction, urging the market maker to reduce the skew in order to avoid revealing inventory information to skew sniffers. The second effect can dominate, for instance, in tiers with a large proportion of skew sniffers, especially when these clients trade infrequently but are highly sensitive to revealed information.
The case of adverse selection once again depends on the relative shape of the $\zeta$ and intensity functions. When adverse selection depends only mildly on the quotes, the second component affects the quote offsets according to the sign of $q+\Delta^{k}$ for bid transactions of size $\Delta^{k}$ and $-q+\Delta^{k}$ for ask transactions of the same size. In such cases, the $\Delta^{k}$ terms contribute to widening the spread, while the $q$ terms encourage additional skewing – opposite to the incentives from the first component. The reasoning is reversed when adverse selection primarily occurs at seemingly unfavorable prices, that is, when informed clients exploit short-term mispricings. However, the second component still pushes in the opposite direction to the first. In both scenarios, the second component may dominate the first in tiers with highly informed clients, especially when trades have a strong impact on prices. However, unlike in the case of price reading, liquidity – through the amplitude of $\Lambda^{n,k}$ – does not play a clear role in this mechanism.
@@ 5.2 Price Reading
Let us consider the case of price reading only. [^11] We consider price reading and adverse selection separately in order to disentangle their respective effects. This separation is without loss of generality in our framework, which is restricted to first-order effects. We have
@@
As discussed earlier, the first term captures the global response to the price reading risk, for trades of size $\Delta^{k}$, uniformly across all tiers, while the second term is tier-specific. This second term is non-linear in inventory and is highly sensitive to the presence and nature of skew sniffers in the tier: a market maker can tolerate skew sniffers with high trading activity because they contribute to global inventory risk reduction, but it is meaningless to show prices to skew sniffers who do not trade (indeed, the term diverges as $\Lambda_{0}^{n,k}\to 0$).
@@
As an illustration, consider a standard size ladder of $(1,2,5,10,20,50)$ million notional and two tiers of equal baseline intensity, with $\Lambda_{0}^{n,k}=(1000,400,300,200,60,40)$ (in day-1 units) and $\kappa^{n,k}=3$ bp-1. This corresponds to a relatively liquid currency pair with a daily turnover exceeding 6 billion notional and a top-of-book spread of approximately $0.7$ bp, assuming a daily volatility of $100$ bp. In the first tier, we assume no price reading, i.e. $w^{1,k}=0$, whereas in the second tier we assume the presence of skew sniffers with $w^{2,k}=1$.
@@
Figure 1: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$ for the two tiers (we illustrate the first and the fourth sizes). Risk aversion $\gamma=10^{-5}$ bp-1 M-1. Other parameters are defined in the text.
@@
Figure 2: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$ for the two tiers (we illustrate the first and the fourth sizes). Risk aversion $\gamma=10^{-5}$ bp-1 M-1. $\Lambda_{0}^{1,k}=(1500,600,450,300,90,60)$, $\Lambda_{0}^{2,k}=(500,200,150,100,30,20)$. Other parameters are defined in the text.
As shown in Figure 1, [^12] We only plot bid quote adjustments because ask quote adjustments are symmetric. optimal pricing for the first tier reflects the global increase in informational risk, i.e. inventory skew is increased. The situation is reversed for the second tier (except of course around the point corresponding to zero inventory), since the skew is reduced through contributions from both the defensive and offensive sides, as the market maker aims to limit information leakage. The effect is particularly pronounced for top-of-book quotes. However, since skew sniffers still contribute significantly to risk reduction through their high $\Lambda_{0}$, the effect diminishes for larger sizes. A decrease in the trading intensity of the second tier leads to a sharp reduction of skew by the market maker across the entire pricing ladder for the second tier with skew sniffers, as illustrated in Figure 2.131313It is worth noting that for larger values of the risk aversion parameter $\gamma$, the market maker tends to reduce the skew shown to skew sniffers through a more important change in the quote on the defensive side than on the offensive side (figure not shown). This is the consequence of the model chosen where changes in price linked to price reading are proportional to skew.

### v2

@@ 2.1 Modeling Framework and Notation
  • Adverse selection. For each $(n,k)\in\mathcal{N}\times\mathcal{K}$, the function $\tilde{\zeta}^{n,k}:\mathbb{R}\to\mathbb{R}_{+}$ is non-decreasing, and satisfies $\lim_{\delta\to-\infty}\tilde{\zeta}^{n,k}(\delta)=0.$ This function captures the impact of informed trading on prices. It is tier-specific (and possibly size-specific), allowing the market maker to account for varying levels of expected informativeness across clients (and transaction sizes). In particular, when $\tilde{\zeta}^{n,k}$ is a rapidly increasing function, executions at prices seemingly unfavorable to the client, from the market maker’s perspective, can have a strong impact on the reference price.
@@ 3.3.3 Optimal Quotes
$\displaystyle=$ $\displaystyle\delta^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}(\delta^{n,k,b*}(q))}\Bigg(D^{k}_{+}f(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}$
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{(q+\Delta^{k})\left(\Lambda^{n,k,b}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,b*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\delta^{n,k,b*}(q)\right)}\Bigg)+o(\varepsilon)$
@@
$\displaystyle\mathfrak{d}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\delta^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}(\delta^{n,k,a*}(q))}\Bigg(D^{k}_{-}f(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\delta^{n,j,a*}(q)-\delta^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}$
$\displaystyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{(q-\Delta^{k})\left(\Lambda^{n,k,a}\zeta^{n,k}\right)^{\prime}\!\left(\delta^{n,k,a*}(q)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\delta^{n,k,a*}(q)\right)}\Bigg)+o(\varepsilon)$
These expressions highlight that the influence of informational risk on optimal quotes can be decomposed into two components. The first component arises from the terms $D^{k}_{\pm}f$, which are finite differences of the function $f$. By definition, these terms reflect the marginal impact of informational risk on the risk-adjusted expected PnL, assuming the market maker does not update their quotes. As such, they aggregate the effects of informational risk across all tiers and are identical for all $n$. The second component is represented by the terms $g^{n,k,b/a}$, which quantify the impact of informational risk on the quotes derived from the baseline value function in the absence of informational risk. Unlike the $D^{k}_{\pm}f$ terms, the $g^{n,k,b/a}$ terms are specific to the considered tier.
@@ 4 Quadratic Hamiltonian Approximation
In the algorithmic market making literature, closed-form approximations of the value function and optimal quotes are frequently used. One approach (see [6]) consists in replacing the HJB equation associated with the control problem by an alternative HJB equation in which the Hamiltonian functions are approximated to second order. The resulting modified HJB equation admits a solution that is quadratic in the inventory variable, with coefficients computable through a Riccati equation and a linear equation. [^8] While we focus on the single-asset case in this article, it is worth emphasizing that this methodology extends naturally to multi-asset market making, thereby offering an effective remedy to the curse of dimensionality.
@@ 4.2 HJB Equation and Value Function
$\left\{\begin{aligned} A_{0}&=\frac{-\rho+\sqrt{\rho^{2}+4\gamma\sigma^{2}F_{+}^{2,1}}}{4F_{+}^{2,1}}\\ B_{0}&=-\frac{2A_{0}}{\rho+2A_{0}F_{+}^{2,1}}\left(A_{0}F_{-}^{2,2}+D_{-}^{1,1}\right)\\ A_{0}^{\prime}&=-\frac{2A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\\ B_{0}^{\prime}&=-\frac{1}{\rho+2A_{0}F_{+}^{2,1}}\left(B_{0}\Sigma^{1,0}_{+}+A_{0}\Sigma^{3,1}_{-}+S-\frac{4A_{0}\Sigma_{+}^{1,0}}{\rho+4A_{0}F_{+}^{2,1}}\left(B_{0}F^{2,1}_{+}+2A_{0}F^{2,2}_{-}+F^{1,1}_{-}\right)\right).\end{aligned}\right$.
@@ 4.3 Optimal Quotes
If the quotes in feedback form associated with $\widehat{\vartheta}_{0}$, i.e. in the case of no informational risk under the quadratic approximation, are denoted by
$\widehat{\delta}^{n,k,b*}(q):=\tilde{\delta}^{n,k,b*}\left(D^{k}_{+}\widehat{\vartheta}_{0}(q)\right)\quad\text{and}\quad\widehat{\delta}^{n,k,a*}(q):=\tilde{\delta}^{n,k,a*}\left(D^{k}_{-}\widehat{\vartheta}_{0}(q)\right),$ (11)
@@
$\displaystyle\widehat{\mathfrak{d}}^{n,k,b*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,b*}(q)+\frac{\varepsilon}{c^{n,k,b}\left(\widehat{\delta}^{n,k,b*}(q)\right)}\left(D^{k}_{+}\widehat{f}(q)+\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,b\,\prime}\!\left(\widehat{\delta}^{n,k,b*}(q)\right)}\right$. (12)
@@
$\displaystyle\widehat{\mathfrak{d}}^{n,k,a*}(q)$ $\displaystyle=$ $\displaystyle\widehat{\delta}^{n,k,a*}(q)+\frac{\varepsilon}{c^{n,k,a}\left(\widehat{\delta}^{n,k,a*}(q)\right)}\left(D^{k}_{-}\widehat{f}(q)-\frac{qw^{n,k}J^{n\,\prime}\!\left(\sum_{j\in\mathcal{K}}w^{n,j}\left(\widehat{\delta}^{n,j,a*}(q)-\widehat{\delta}^{n,j,b*}(q)\right)\right)}{{\Delta^{k}}\Lambda^{n,k,a\,\prime}\!\left(\widehat{\delta}^{n,k,a*}(q)\right)}\right$. (13)
@@ 4.4 Analysis in the Symmetric Case
Given that $\widehat{f}$ is available in closed form, the above expressions are easier to handle than those in Section 3.3.3. This allows us to examine the underlying mechanisms in detail. For simplicity, we focus on a symmetric setting with identical bid- and ask-side functions, in the limiting case where $\rho\to 0$.
In this case indeed, we have $B_{\varepsilon}=B^{\prime}_{\varepsilon}=0$, and the first-order adjustment to the value function is, up to an additive constant, given by
$\widehat{f}(q)=\frac{1}{2}\frac{\Sigma_{+}^{1,0}}{F_{+}^{2,1}}\,q^{2},$
where [^10] We removed the $b$ and $a$ superscripts because of our assumption of symmetry.
@@
Let us first consider a tier where clients trade on strong signals and would execute in the direction of their signal at virtually any reasonable price given prevailing market conditions. Such behavior is plausible for clients acting on mid- or long-term signals. In this case, the slopes of the corresponding $\zeta$ functions are insufficiently steep to compensate the decrease of intensity functions, resulting in a reduction in the concavity of the value function. The mechanism is as follows: if the market maker holds a positive (resp. negative) inventory, then by skewing – through a decrease (resp. increase) in both bid and ask prices – they raise the absolute and relative probability of executing a sell (resp. buy) order that moves the inventory closer to zero and therefore increase the likelihood that adverse selection – whose dependence on quotes is mild – induces a price increase (resp. decrease), thereby increasing the value of the remaining inventory. Although adverse selection reduces the constant term in the approximation of the value function, this mechanism tends to flatten its shape by lowering its concavity.
@@
The second component of the optimal quotes, which is tier-specific and relates to the local adjustments of quotes, consists of two terms: one related to price reading and the other to adverse selection.
@@
Taken together, the two components of the impact of price reading on optimal quotes lead to a subtle trade-off. When the inventory is zero, the second component vanishes and only the first component remains: a widening of the bid-ask spread to compensate for informational risk. When the inventory is nonzero, the first component encourages stronger skewing to reduce the time spent in unbalanced positions, while the second component pushes in the opposite direction, urging the market maker to reduce the skew in order to avoid revealing inventory information to skew sniffers. The second effect dominates for instance in tiers with a large proportion of skew sniffers that are highly sensitive to revealed information but trade infrequently.
The case of adverse selection once again depends on the relative shape of the $\zeta$ and $\Lambda$ functions. When adverse selection depends only mildly on the quotes, the second component affects the quote offsets according to the sign of $q+\Delta^{k}$ for bid transactions of size $\Delta^{k}$ and $-q+\Delta^{k}$ for ask transactions of the same size. In such cases, the $\Delta^{k}$ terms contribute to widening the spread, while the $q$ terms encourage additional skewing – opposite to the incentives from the first component. The reasoning is reversed when adverse selection primarily occurs at seemingly unfavorable prices, that is, when informed clients exploit short-term mispricings. However, the second component still pushes in the opposite direction to the first. In both scenarios, the second component may dominate the first in tiers with highly informed clients, especially when trades have a strong impact on prices. However, unlike in the case of price reading, liquidity – through the amplitude of $\Lambda^{n,k}$ – does not play a clear role in this mechanism.
@@ 5.2 Price Reading
Let us first consider the case of price reading alone. [^11] We consider price reading and adverse selection separately in order to disentangle their respective effects. This separation is without loss of generality in our framework, which is restricted to first-order effects and therefore ignores any cross-effects. We have
@@
As discussed earlier, the first term captures the global response to the price reading risk for trades of size $\Delta^{k}$, uniformly across all tiers, while the second term is tier-specific. This second term is non-linear in inventory and highly sensitive to the presence and nature of skew sniffers in the tier: market makers need to only slightly reduce the skew shown to skew sniffers with high trading activity (as they contribute to global inventory risk reduction), whereas they must reduce it more substantially for those who trade infrequently.
@@
As an illustration, consider a standard size ladder of $(1,2,5,10,20,50)$ million (hereafter M) notional and two tiers with total baseline intensities $\Lambda_{0}^{1,k}+\Lambda_{0}^{2,k}=(2000,800,600,400,120,80)$ day-1 and $\kappa^{n,k}=3$ bp-1. This corresponds to a relatively liquid currency pair with a daily turnover exceeding $6$ billion notional and a top-of-book spread of approximately $0.7$ bp. We also assume a daily volatility of $100$ bp and a market maker’s risk aversion parameter of $\gamma=10^{-4}$ bp-1 M-1.
In the first tier, we assume no price reading, i.e., $w^{1,k}=0$, whereas in the second tier we assume the presence of skew sniffers with $w^{2,k}=\frac{\exp(-\sqrt{\Delta^{k}})}{1000}$. We define the safe volume share (SVS) as the ratio $\text{SVS}=\frac{\Lambda_{0}^{1,k}}{\Lambda_{0}^{1,k}+\Lambda_{0}^{2,k}}$, which represents the percentage of trading intensity not subject to price reading.
@@
Figure 1: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$, with the second tier subject to price reading (we illustrate the first and the fourth sizes) – $\text{SVS}=25\%$.
@@
Figure 2: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$, with the second tier subject to price reading (we illustrate the first and the fourth sizes) – $\text{SVS}=75\%$.
As shown in Figures 1 and 2, [^12] We only plot bid quote adjustments because ask quote adjustments are symmetric. optimal pricing for the first tier reflects the global increase in informational risk: spread widening (more visible for large sizes) and an increase in skew. For the second tier, the additional skew related to the global increase in informational risk is counterbalanced by the objective to reduce information leakage, which implies a reduction in the skew shown to skew sniffers. In both figures, we clearly see that the skews are indeed lower for the second tier than for the first. Moreover, the de-skewing effect is particularly pronounced for top-of-book quotes in both the low and high SVS cases: the skew associated with top-of-book quotes for tier two is not only lower than for tier one, but it is in fact lower than in the absence of skew sniffers. The case of larger sizes is notable as it depends strongly on the value of SVS. When $\text{SVS}=25\%$ (see Figure 1), the skew shown to clients of tier two for large sizes is still larger than in the absence of skew sniffers because they contribute significantly to risk reduction through their relatively high trading intensity. In contrast, when $\text{SVS}=75\%$ (see Figure 2), the de-skewing effect dominates even for large sizes. In all cases, it is noteworthy that the contribution of the defensive side to de-skewing is always higher than that of the offensive side.
Let us now simulate the PnL of a market maker facing price reading as described above, over a time window $[0,T]$ where $T=10^{4}$ s, for different values of the safe volume share SVS and several quoting strategies: (i) the “No Action” strategy, which does not take price reading into account and uses the quotes from Eq. (11), (ii) the “No Skew” strategy, which consists of withdrawing skew from prices shown to the tier of price readers, and (iii) the “Optimal” strategy, which uses the first-order approximations of Eqs. (12) and (13) with $\varepsilon=1$. [^13] We added a floor at $0$ for $\widehat{\mathfrak{d}}^{n,k,b/a*}(q)$ to avoid excessive aggressiveness on the risk-increasing side, which may arise in extreme situations where the applicability of the first-order approximations is questionable.
Figures 3, 4 and 5 illustrate the average PnL, its standard deviation, and the ratio between the two (as a risk-adjusted performance measure), respectively, over $10^{5}$ independent simulations for several values of SVS when using the above three strategies. The “No Reading” baseline corresponds to the absence of skew sniffers, i.e., $w^{2,k}=0$, and uses the quoting strategy of Eq. (11). We clearly see that the presence of skew sniffers implies a decrease in the expected PnL compared to the baseline if quotes are not modified to account for price reading (“No Action” strategy). However, the risk is not affected, and no dependence on the safe volume share is visible. The naïve “No Skew” approach (consisting of showing no skew to the skew sniffers) helps restore the PnL but leads to a considerable increase in risk, particularly if the price readers trade a lot (i.e., for low values of SVS). The pricing strategy based on our first-order approximations not only increases the PnL compared to that of the “No Action” strategy but also reduces the risk. This strategy is more efficient when the fraction of skew sniffers is low. [^14] Note that a PnL higher than the baseline is not unexpected, as the strategy effectively introduces a controlled price drift. The risk-adjusted performance measures of Figure 5 clearly emphasize the benefits of our first-order approximations.
Figure 3: Average values of the simulated market maker’s PnL as a function of safe volume share for the different strategies.
Figure 4: Values of the standard deviation for the simulated market maker’s PnL as a function of safe volume share for the different strategies.
Figure 5: Risk-adjusted performance measure associated with the simulated market maker’s PnL as a function of safe volume share for the different strategies.